=== 2019/5/29(Wed.) ===

==== Schedule ====
 * --(5/22-23: MZ release from PSL)--
 * --(5/20-24: Exchanging cryo valves.)--
 * 5/27-6/2: Y arm mass lock.
 * 5/29: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (specially BS)
 * 6/3-6/14: Installation of TMS for ETMY
 * 6/8: Engineering run

 * ITMX
  * --(5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm)--
  * --(5/22AM~: tuned off shield cooler, finesse measurement for Xarm)--

 * ETMX
  * --(5/2: Cooling start)--

 * ITMY
  * --(5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm)--

 * ETMY
  * --(5/17~: tuned off ETMY cryostat cooler)--

 * Long term:
  * 4/22-5/31: Yarm commissioning	
  * 5/27-6/30: DRMI commissioning
  * 7/1-8/9: DRFPMI commissioning
  * 8/12-10/4: DRFPMI ASC
  * 10/7-11/15: Post Commissioning
  * 11/18-12/13: Noise hunting


==== Note ====

==== Night work ====

==== Weekend ====

==== Report of work ====
 * COM: GR for X arm was locked again. The problem was that the shutter of GR was half closed since last Wednesday.
  * Some resonances of 3.6H and 5.1Hz were excited when length control signal was injected to hierarchy control point.
 * CRY: ITMY reached 29K.
 * TYPE A: Measurement of TF for Type A is going on from this morning.

==== Work plan ====
 * DGS: Maintenance for replacing k1lsc0.
 * COM: Finesse measurement for y arm with low temperature. Contamination can happen with low temperature.
 * Type B: Smooth watch dog will be implemented into some Type Bs in the maintenance time.
 * IOO: Laser safety button will be available this afternoon.
 * VIS: Status monitor for VIS will show up in the control room.